Data for daily returns is provided below for Chinese Stocks. Calculate expected returns and standard deviation and recommend which counter you would invest in based on expected returns and secondly on risk.

Listed Company Code_Comcd

Stock Code_Stkcd

Latest Stock Name_Lstknm

Date_Date

Daily Return_Dret

C603508

603508

Thinker

2018-01-18

0,0138

C603508

603508

Thinker

2018-01-19

-0,0054

C603508

603508

Thinker

2018-01-22

0,0116

C603508

603508

Thinker

2018-01-23

0,0018

C603508

603508

Thinker

2018-01-24

0,014

C603508

603508

Thinker

2018-01-25

-0,0123

C603508

603508

Thinker

2018-01-26

-0,0081

C603508

603508

Thinker

2018-01-29

-0,0154

C603508

603508

Thinker

2018-01-30

-0,0508

C603508

603508

Thinker

2018-01-31

-0,0634

C603508

603508

Thinker

2018-02-01

-0,0129

C603508

603508

Thinker

2018-02-02

0,0457

C603508

603508

Thinker

2018-02-05

-0,0281

C603508

603508

Thinker

2018-02-06

-0,0315

C603508

603508

Thinker

2018-02-07

0,0166

C603508

603508

Thinker

2018-02-08

0,0122

C603508

603508

Thinker

2018-02-09

-0,019

C603508

603508

Thinker

2018-02-12

0,0182

C603508

603508

Thinker

2018-02-13

0,0009

C603508

603508

Thinker

2018-02-14

0,0026

C603508

603508

Thinker

2018-02-22

0,0102

C603508

603508

Thinker

2018-02-23

0,0017

C603508

603508

Thinker

2018-02-26

0,0254

C603508

603508

Thinker

2018-02-27

0,0245
Answer #1
R D=R-(-0.0019875) E=D^2
Date Daily Return Deviation from Expected Deviation Squared
1/18/2018 0.0138 0.0157875 0.000249245
1/19/2018 -0.0054 -0.0034125 1.16452E-05
1/22/2018 0.0116 0.0135875 0.00018462
1/23/2018 0.0018 0.0037875 1.43452E-05
1/24/2018 0.014 0.0159875 0.0002556
1/25/2018 -0.0123 -0.0103125 0.000106348
1/26/2018 -0.0081 -0.0061125 3.73627E-05
1/29/2018 -0.0154 -0.0134125 0.000179895
1/30/2018 -0.0508 -0.0488125 0.00238266
1/31/2018 -0.0634 -0.0614125 0.003771495
2/1/2018 -0.0129 -0.0109125 0.000119083
2/2/2018 0.0457 0.0476875 0.002274098
2/5/2018 -0.0281 -0.0261125 0.000681863
2/6/2018 -0.0315 -0.0295125 0.000870988
2/7/2018 0.0166 0.0185875 0.000345495
2/8/2018 0.0122 0.0141875 0.000201285
2/9/2018 -0.019 -0.0170125 0.000289425
2/12/2018 0.0182 0.0201875 0.000407535
2/13/2018 0.0009 0.0028875 8.33766E-06
2/14/2018 0.0026 0.0045875 2.10452E-05
2/22/2018 0.0102 0.0121875 0.000148535
2/23/2018 0.0017 0.0036875 1.35977E-05
2/26/2018 0.0254 0.0273875 0.000750075
2/27/2018 0.0245 0.0264875 0.000701588
Total -0.0477 SUM 0.014026166
A Sum of Returns -0.0477
B Number of observations                          24
C=A/B Expected Return        (0.0019875)
X Sum of Deviation squared        0.01402617
Y=X/(B-1) Variance of return          0.0006098
Z=Square Root(Y) Standard Deviation of Return 0.024694803
EXPECTED RETURN          (0.001988) -0.1988%
STANDARD DEVIATION             0.024695 2.4695%